OpenStudy (anonymous):

what is the mass and weight of a 10kg object on earth

OpenStudy (anonymous):

Mass is the amount of matter an object is composed of. 10 kg IS a mass. The weight of an object is the mass of the object times the rate at which it accelerates under gravity. This is the force acting on the object. On Earth, 'g' is 9.8 m/s^2. W = 10 * 9.8 = 980 N